### v1

@@ 3.1 Proper orthogonal decomposition
$\Xi_{\Delta}\coloneqq\{\Psi(\omega^{n}(t_{k}),p\Delta t),\;\forall p\in\{0,...,N_{\Theta}\},\,\omega^{n}(t_{k})\in E_{k|r}\},$
where $E_{k|r}\coloneqq\{\omega^{n}(t_{k})\sim\pi(\omega(t_{k})|\mathbf{d}_{r})\}_{n=1}^{N_{\Xi}}$, $N_{\Xi}\in\mathbb{N}^{+}$, is an ensemble of random initial states, and $\Delta t\coloneqq T/N_{\Theta}$, $N_{\Theta}\in\mathbb{N}^{+}$, is a time step, sufficiently small to capture the system’s dynamics. Quasi-optimal approximation spaces can be constructed by minimizing the cost function
$J_{\Delta}({\mathcal{V}}_{\varepsilon})\coloneqq\dfrac{1}{N_{\Xi}(N_{\Theta}+1)}\sum_{n=1}^{N_{\Xi}}\sum_{p=0}^{N_{\Theta}}\|(\text{Id}-\Pi_{{\mathcal{V}}_{\varepsilon}})\omega^{n}(t_{k}+p\Delta t)\|_{\mathcal{V}}^{2},$
with $\omega^{n}(t_{k}+p\Delta t)\in\Xi_{\Delta}$. The solution of the resulting optimization problem is closely related to the eigenvalue problem associated with the linear functional $\mathcal{R}:\mathcal{V}\rightarrow\mathcal{V}$, defined as
$\mathcal{R}\coloneqq\dfrac{1}{N_{\Xi}(N_{\Theta}+1)}\sum_{n=1}^{N_{\Xi}}\sum_{p=0}^{N_{\Theta}}\omega^{n}(t_{k}+p\Delta t)(\omega^{n}(t_{k}+p\Delta t),\,\cdot\,)_{\mathcal{V}}.$
@@
Leveraging the result expressed in Equation (7), we can efficiently identify the optimal size ensuring an approximation error below a specified threshold, i.e., for a given $\varepsilon_{k}\in\mathbb{R}^{+}$, we can select the smallest $N_{\varepsilon}$ for which $\min J_{\Delta}(\mathcal{V}_{\varepsilon})\leq\varepsilon_{k}$. However, determining an appropriate value for $\varepsilon_{r}$ remains an open problem. In the rest of this section, we propose a tolerance selection strategy linked to the covariance of the distribution characterizing the uncertainty in the system’s state. Specifically, we introduce a novel method for dynamically choosing $\varepsilon$, relying on the average distance between two states independently sampled from the prior distribution $\pi(\omega(t_{k}),\mathbf{d}_{k})$. More in details, let $\omega(t_{k})$ and ${\omega^{\prime}}(t_{k})$, be two statistically independent random variables distributed according to the same probability law $\pi(\omega(t_{k})|\mathbf{d}_{k})$, and let $\varepsilon_{r}\in(0,1)$ be a user-defined relative tolerance for the projection error, then we define the local tolerance as
$\varepsilon_{k}\coloneqq$ $\;\varepsilon_{r}\int_{\mathcal{V}}\int_{\mathcal{V}}\|\omega(t_{k})-{\omega^{\prime}}(t_{k})\|_{\mathcal{V}}^{2}\,\pi(d{\omega^{\prime}}(t_{k})|\mathbf{d}_{k})\,\pi(d\omega(t_{k})|\mathbf{d}_{k})$ (8)
$=$  $\;2\varepsilon_{r}\int_{\mathcal{V}}\|\omega(t_{k})-\overline{\omega}(t_{k})\|_{\mathcal{V}}^{2}\,\pi(d\omega(t_{k})|\mathbf{d}_{k})=2\varepsilon_{r}\,\text{tr}\left(\mathbf{C}^{\omega,\omega}_{k}\,\cdot\,,\,\cdot\,\right)_{\mathcal{V}},$
where $\mathbf{C}^{\omega,\omega}_{k}:\mathcal{V}\rightarrow\mathcal{V}$ denotes the covariance of $\omega(t_{k})\sim\pi(\omega(t_{k}),\mathbf{d}_{k})$, and $\overline{\omega}(t_{k})\in\mathcal{V}$ its mean. If the exact probability measure is approximated by an empirical measure, as it happens in the EnKF algorithms, the covariance can be approximated by an appropriate Monte Carlo estimator, as shown in Section 4.3.
@@ 4 Adaptive hierarchical ensemble Kalman filter with reduced basis models
In this work, we propose to generate these models via adaptive model order reduction of the full order model (1). This operation is conducted dynamically, employing a limited number of costly expensive training data obtained from the full order model alongside several inexpensive data from surrogate models trained in previous steps. With respect to the use of training data from the full order model alone, this allows for better surrogate models that can retain information and yield more precise reconstructions. For the sake of simplicity, we focus on a two-level setting, and we refer to Section 4.4 for a brief description of the generalization to an arbitrary number of levels.
@@
In the next sections we describe the analysis step of the ML-EnKF Section (4.1) and of the MF-EnKF (Section 4.2), while in Section 4.3 we propose a novel strategy for the prediction step.
@@ 4.1 The ML-EnKF analysis step
In contrast with the standard EnKF, the analysis step of the ML-EnKF updates the states in the ensembles with the empirical Kalman gain $\mathbf{K}_{k}^{\text{ML}}\coloneqq\mathbf{Q}^{\text{ML}}_{k}(\mathbf{P}^{\text{ML}}_{k}+\bm{\Sigma})^{-1}$ where the state-measurement covariance $\mathbf{Q}^{\text{ML}}_{k}:\mathcal{Y}\rightarrow\mathcal{V}$ and the measurement-measurement covariance $\mathbf{P}^{\text{ML}}_{k}:\mathcal{Y}\rightarrow\mathcal{Y}$ are obtained via multi-level Monte Carlo estimation. This operation involves (1) the telescopic expansion of the two covariances in terms of partial covariances associated with models of different accuracy and ensembles of different sizes, and (2) a regularization step to ensure the stability of the method. Considering only the three ensembles introduced in the previous section, and two levels of accuracy associated with the approximation spaces $\mathcal{V}$ and $\mathcal{V}_{\varepsilon}^{k}$, the unregularized covariances $\widetilde{\mathbf{Q}}^{\text{ML}}_{k}:\mathcal{Y}\rightarrow\mathcal{V}$ and $\widetilde{\mathbf{P}}^{\text{ML}}_{k}:\mathcal{Y}\rightarrow\mathcal{Y}$ can be defined as
$\widetilde{\mathbf{Q}}^{\text{ML}}_{k}$ $\coloneqq\mathbf{Q}^{P,P}_{k}-\mathbf{Q}^{C,C}_{k}+\mathbf{Q}^{A,A}_{k},$ (9)
$\widetilde{\mathbf{P}}^{\text{ML}}_{k}$ $\coloneqq\mathbf{P}^{P,P}_{k}-\mathbf{P}^{C,C}_{k}+\mathbf{P}^{A,A}_{k}.$
@@
Although, by construction, the covariance matrices are symmetric and positive semi-definite, their combination in (9) may yield an indefinite symmetric covariance $\widetilde{\mathbf{P}}^{\text{ML}}_{k}$. The presence of a negative part of the spectrum has the potential to trigger instabilities in the data assimilation process. To counteract this issue, regularization techniques are usually employed anderson2012 ; bickel2008_1 ; bickel2008 ; chernov2021 ; hoel2016 ; olivier2004 . In this work, we propose a regularization based on the projection of both $\widetilde{\mathbf{P}}^{\text{ML}}_{k}$ and $\widetilde{\mathbf{Q}}^{\text{ML}}_{k}$ to the cone of positive semi-definite linear operators, so as to move the unstable measurement subspace to their kernel. More precisely, let $(\lambda_{i},\mathbf{p}_{i})\in\mathbb{R}\times\mathcal{Y}$, for all $i\in\{1,...,N_{D}\}$, be the eigenpairs of $\widetilde{\mathbf{P}}^{\text{ML}}_{k}$. The regularized covariances $\mathbf{P}^{\text{ML}}_{k}:\mathcal{Y}\rightarrow\mathcal{Y}$ and $\mathbf{Q}^{\text{ML}}_{k}:\mathcal{Y}\rightarrow\mathcal{V}$ are computed as follows:
$\mathbf{Q}^{\text{ML}}_{k}$ $\coloneqq\sum_{\begin{subarray}{c}{i}=1\\ \lambda_{i}\geq 0\end{subarray}}^{N_{{D}}}\widetilde{\mathbf{Q}}^{\text{ML}}_{k}\mathbf{p}_{i}(\mathbf{p}_{i},\,\cdot\,)_{\mathcal{Y}},$
$\mathbf{P}^{\text{ML}}_{k}$ $\coloneqq\sum_{\begin{subarray}{c}{i}=1\\ \lambda_{i}\geq 0\end{subarray}}^{N_{{D}}}\widetilde{\mathbf{P}}^{\text{ML}}_{k}\mathbf{p}_{i}(\mathbf{p}_{i},\,\cdot\,)_{\mathcal{Y}}=\sum_{\begin{subarray}{c}{i}=1\\ \lambda_{i}\geq 0\end{subarray}}^{N_{{D}}}\lambda_{{i}}\mathbf{p}_{i}(\mathbf{p}_{i},\,\cdot\,)_{\mathcal{Y}}.$
@@
Finally, the regularized covariances are used to define the multi-level Kalman gain $\mathbf{K}_{k}^{\text{ML}}\coloneqq\mathbf{Q}^{\text{ML}}_{k}(\mathbf{P}^{\text{ML}}_{k}+\bm{\Sigma})^{-1}$, which is employed to update the states of the three ensembles. Specifically, for all $n\in\{1,...,N_{P}\}$ and $m\in\{1,...,N_{A}\}$, we update the ensemble members as
@@ 4.2 The MF-EnKF analysis step
$\mathbf{P}^{\text{MF}}_{k}$ $=\mathbf{P}^{P,P}_{k}+\dfrac{1}{4}\left(\mathbf{P}^{C,C}_{k}+\mathbf{P}^{A,A}_{k}\right)-\dfrac{1}{2}\left(\mathbf{P}^{P,C}_{k}+\mathbf{P}^{C,P}_{k}\right).$
@@
The multi-fidelity Kalman gain is then defined as $\mathbf{K}_{k}^{\text{MF}}\coloneqq\mathbf{Q}^{\text{MF}}_{k}(\mathbf{P}^{\text{MF}}_{k}+\tfrac{1}{2}\bm{\Sigma})^{-1}$, with $\mathbf{P}^{\text{MF}}_{k}$ and $\mathbf{Q}^{\text{MF}}_{k}$ as in (12). The states of the three ensembles are updated, for all $n\in\{1,...,N_{P}\}$ and $m\in\{1,...,N_{A}\}$, as
@@
$\overline{\omega}_{k|k}^{\text{MF}}\coloneqq\overline{\omega}_{k|k}^{P}+\dfrac{1}{2}(\overline{\omega}_{k|k}^{A}-\overline{\omega}_{k|k}^{{C}}),$
where
$\overline{\omega}_{k|k}^{\Diamond}=\dfrac{1}{N_{\Diamond}}\textstyle\sum_{n=1}^{N_{\Diamond}}\tilde{\omega}_{k|k}^{{\Diamond,n}},\quad\Diamond\in\{P,C,A\}.$
@@ 4.3 The prediction step: an inflation-deflation adaptive reduced basis approach
At each iteration of the EnKF algorithm, we construct two subspaces of $\mathcal{V}$: the first one $\mathcal{V}_{\varepsilon}^{k}$, called the inflated space, is the reduced basis space; the second one $\mathcal{W}_{\varepsilon}^{k}$, called the deflated space, is the approximation space that transfers information from one assimilation window to the subsequent one. Before starting the data assimilation process, along with the initial ensembles, it is necessary to specify the starting deflated space, $\mathcal{W}_{\varepsilon}^{0}$, to incorporate any potential prior knowledge regarding the model’s reducibility. In the absence of more information, one can assume $\mathcal{W}_{\varepsilon}^{0}\equiv\mathcal{V}$. Next, at the $k$th iteration of the data assimilation algorithm, the prediction step starts with the forecast of the post-analysis principal ensemble, $E_{k|k}^{P}$, over the assimilation window $\mathcal{I}_{k}$. Specifically, the high-fidelity model $\Psi(\,\cdot\,,t_{k+1}-t_{k})$ is employed to update the states in the principal ensemble
$\omega_{k+1|k}^{P,n}=\Psi(\omega_{k|k}^{P,n},t_{k+1}-t_{k}),\quad\forall\,n\in\{1,...,N_{P}\}.$ (15)
@@
$\mathcal{V}_{\varepsilon}^{k+1}\coloneqq\text{span}\{\text{POD}(\Xi_{\Delta,k}^{\text{HF}},\varepsilon^{\Diamond}_{k}/2),\mathcal{W}_{\varepsilon}^{k}\},$
@@
$\varepsilon^{\text{MF}}_{k}=\dfrac{1}{2}\dfrac{\varepsilon_{r}}{N_{A}-1}\sum_{n=1}^{N_{A}}\|\omega_{k|k}^{A,m}-\overline{\omega}_{k|k}^{A}\|_{\mathcal{V}}^{2}+\dfrac{2\varepsilon_{r}}{N_{P}-1}\sum_{n=1}^{N_{P}}\Big{(}\|\omega_{k|k}^{P,n}-\overline{\omega}_{k|k}^{P}\|_{\mathcal{V}}^{2}$ (17)
$-(\omega_{k|k}^{P,n}-\overline{\omega}_{k|k}^{P},\omega_{k|k}^{C,n}-\overline{\omega}_{k|k}^{C})_{\mathcal{V}}+\dfrac{1}{4}\|\omega_{k|k}^{C,n}-\overline{\omega}_{k|k}^{C}\|_{\mathcal{V}}^{2}\Big{)}.$
@@
The resulting approximation space $\mathcal{V}_{\varepsilon}^{k+1}$ of size $N_{\varepsilon}^{k+1}$ is used to construct the surrogate model $\Psi^{k+1}_{\varepsilon}(\,\cdot\,,t_{k+1}-t_{k})$, following the procedure of Section 3. The updated model is then employed to advance the states in the control and ancillary ensembles, namely
$\omega_{k+1|k}^{C,n}$ $=\Psi^{k+1}_{\varepsilon}(\Pi_{\mathcal{V}_{\varepsilon}^{k+1}}\omega_{k|k}^{P,n},t_{k+1}-t_{k}),$ $\quad\forall\,n\in\{1,...,N_{P}\},$ (18)
$\omega_{k+1|k}^{A,m}$ $=\Psi^{k+1}_{\varepsilon}(\Pi_{\mathcal{V}_{\varepsilon}^{k+1}}\omega_{k|k}^{A,m},t_{k+1}-t_{k}),$ $\quad\forall\,m\in\{1,...,N_{A}\}.$
@@
respectively. These are used to deflate $\mathcal{V}_{\varepsilon}^{k+1}$, retaining only the modes actively employed by the surrogate model: if $\Xi_{\Delta}^{\text{LF}}\coloneqq\mathcal{E}_{k}^{C}\cup\mathcal{E}_{k}^{A}$, the updated deflated space is
$\mathcal{W}_{\varepsilon}^{k+1}\coloneqq\text{POD}(\Xi_{\Delta}^{\text{LF}},\varepsilon^{\Diamond}_{k}/2),$

### v2

@@ 3.1 Proper orthogonal decomposition
$\Xi_{\Delta}:=\{{\color[rgb]{0,0,0}\Psi(\omega^{n}(t_{k}),t_{k},t_{k}+p\Delta t)},\;\forall p\in\{0,...,N_{\Theta}\},\,\omega^{n}(t_{k})\in E_{k|r}\},$
where $E_{k|r}:=\{\omega^{n}(t_{k})\sim\pi(\omega(t_{k})|\mathbf{d}_{r})\}_{n=1}^{N_{\Xi}}$, $N_{\Xi}\in\mathbb{N}^{+}$, is an ensemble of random initial states, and $\Delta t:=T/N_{\Theta}$, $N_{\Theta}\in\mathbb{N}^{+}$, is a time step, sufficiently small to capture the system’s dynamics. Quasi-optimal approximation spaces can be constructed by minimizing the cost function
$J_{\Delta}({\mathcal{V}}_{\varepsilon}):=\dfrac{1}{N_{\Xi}(N_{\Theta}+1)}\sum_{n=1}^{N_{\Xi}}\sum_{p=0}^{N_{\Theta}}\big{\|}(\text{Id}-\Pi_{{\mathcal{V}}_{\varepsilon}})\omega^{n}(t_{k}+p\Delta t)\big{\|}_{\mathcal{V}}^{2},$
with $\omega^{n}(t_{k}+p\Delta t)\in\Xi_{\Delta}$. The solution of the resulting optimization problem is closely related to the eigenvalue problem associated with the linear operator $\mathcal{R}:\mathcal{V}\rightarrow\mathcal{V}$, defined as
$\mathcal{R}:=\dfrac{1}{N_{\Xi}(N_{\Theta}+1)}\sum_{n=1}^{N_{\Xi}}\sum_{p=0}^{N_{\Theta}}\omega^{n}(t_{k}+p\Delta t)(\omega^{n}(t_{k}+p\Delta t),\,\cdot\,)_{\mathcal{V}}.$
@@
Leveraging the result expressed in Equation (7), we can efficiently identify the optimal size ensuring an approximation error below a specified threshold, i.e., for a given $\varepsilon_{k}\in\mathbb{R}^{+}$, we can select the smallest $N_{\varepsilon}$ for which $\min J_{\Delta}(\mathcal{V}_{\varepsilon})\leq\varepsilon_{k}$. However, determining an appropriate value for $\varepsilon_{k}$ remains an open problem. Ideally, the value should be lower than the average distance between states sampled from $\pi(\omega(t_{k})|\mathbf{d}_{r})$ to avoid confusing approximation error and state uncertainty. Based on this intuition, in the following section, we propose a new tolerance selection strategy that links $\varepsilon_{k}$ to the covariance of $\pi(\omega(t_{k})|\mathbf{d}_{r})$.
Let $\omega(t_{k})$ and ${\omega^{\prime}}(t_{k})$, be two statistically independent random variables distributed according to the same probability law $\pi(\omega(t_{k})|\mathbf{d}_{k})$, and let $\varepsilon_{r}\in(0,1)$ be a user-defined relative tolerance for the projection error, then we define the local tolerance as
$\varepsilon_{k}:=$ $\;\varepsilon_{r}\int_{\mathcal{V}}\int_{\mathcal{V}}\big{\|}\omega(t_{k})-{\omega^{\prime}}(t_{k})\big{\|}_{\mathcal{V}}^{2}\,\pi(d{\omega^{\prime}}(t_{k})|\mathbf{d}_{k})\,\pi(d\omega(t_{k})|\mathbf{d}_{k})$ (8)
$=$  $\;2\varepsilon_{r}\int_{\mathcal{V}}\big{\|}\omega(t_{k})-\overline{\omega}(t_{k})\big{\|}_{\mathcal{V}}^{2}\,\pi(d\omega(t_{k})|\mathbf{d}_{k})=2\varepsilon_{r}\,\text{tr}\left(\mathbf{C}^{\omega,\omega}_{k}\,\cdot\,,\,\cdot\,\right)_{\mathcal{V}},$
where $\mathbf{C}^{\omega,\omega}_{k}:\mathcal{V}\rightarrow\mathcal{V}$ denotes the covariance of $\omega(t_{k})\sim\pi(\omega(t_{k}),\mathbf{d}_{k})$, and $\overline{\omega}(t_{k})\in\mathcal{V}$ its mean. If the exact probability measure is approximated by an empirical measure, as it happens in the EnKF algorithms, the covariance can be approximated by an appropriate Monte Carlo estimator, as shown in Subsection 4.3.
@@ 4 Adaptive hierarchical ensemble Kalman filter with reduced basis models
In this work, we propose to generate these models via adaptive reduced order modeling of the full model (1). This operation is conducted dynamically, employing a limited number of costly expensive training data obtained from the full order model alongside several inexpensive data from surrogate models trained in previous steps. With respect to the use of training data from the full order model alone, this allows for better surrogate models that can retain information and yield more precise reconstructions. For the sake of simplicity, we focus on a two-level setting, and we refer to Subsection 4.4 for a brief description of the generalization to an arbitrary number of levels.
@@
In the next sections we describe the analysis step of the ML-EnKF (Subsection 4.1) and of the MF-EnKF (Subsection 4.2), while in Subsection 4.3 we propose a novel strategy for the prediction step.
@@ 4.1 The ML-EnKF analysis step
In contrast with the standard EnKF, the analysis step of the ML-EnKF updates the states in the ensembles with the empirical Kalman gain $\mathbf{K}_{k}^{\text{ML}}:=\mathbf{Q}^{\text{ML}}_{k}(\mathbf{P}^{\text{ML}}_{k}+\bm{\Sigma})^{-1}$ where the state-measurement covariance $\mathbf{Q}^{\text{ML}}_{k}:\mathcal{Y}\rightarrow\mathcal{V}$ and the measurement-measurement covariance $\mathbf{P}^{\text{ML}}_{k}:\mathcal{Y}\rightarrow\mathcal{Y}$ are obtained via multi-level Monte Carlo estimation. This operation involves (1) the telescopic expansion of the two covariances in terms of partial covariances associated with models of different accuracy and ensembles of different sizes, and (2) a regularization step to ensure the stability of the method. Considering only the three ensembles introduced in the previous section, and two levels of accuracy associated with the approximation spaces $\mathcal{V}$ and $\mathcal{V}_{\varepsilon}^{k}$, the unregularized covariances $\widetilde{\mathbf{Q}}^{\text{ML}}_{k}:\mathcal{Y}\rightarrow\mathcal{V}$ and $\widetilde{\mathbf{P}}^{\text{ML}}_{k}:\mathcal{Y}\rightarrow\mathcal{Y}$ can be defined as
$\widetilde{\mathbf{Q}}^{\text{ML}}_{k}$ $:=\mathbf{Q}^{P,P}_{k}-\mathbf{Q}^{C,C}_{k}+\mathbf{Q}^{A,A}_{k},\qquad$ (9)
$\widetilde{\mathbf{P}}^{\text{ML}}_{k}$ $:=\mathbf{P}^{P,P}_{k}-\mathbf{P}^{C,C}_{k}+\mathbf{P}^{A,A}_{k}.$
@@
Although, by construction, the covariance matrices are symmetric and positive semi-definite, their combination in (9) may yield an indefinite symmetric covariance $\widetilde{\mathbf{P}}^{\text{ML}}_{k}$. The presence of a negative part of the spectrum has the potential to trigger instabilities in the data assimilation process. To counteract this issue, regularization techniques are commonly used, such as projecting the matrices on the cone of positive semi-definite matrices via spectrum thresholding bickel2008_1 ; chernov2021 ; hoel2016 , banding or tapering the matrix anderson2012 ; bickel2008 ; olivier2004 , or performing the combination in a log-Euclidean geometry maurais23 .
In this work, we propose a regularization based on the projection of both $\widetilde{\mathbf{P}}^{\text{ML}}_{k}$ and $\widetilde{\mathbf{Q}}^{\text{ML}}_{k}$ to the cone of positive semi-definite linear operators, so as to move the unstable measurement subspace to their kernel. More precisely, let $(\lambda_{i},\mathbf{p}_{i})\in\mathbb{R}\times\mathcal{Y}$, for all $i\in\{1,...,N_{D}\}$, be the eigenpairs of $\widetilde{\mathbf{P}}^{\text{ML}}_{k}$. The regularized covariances $\mathbf{P}^{\text{ML}}_{k}:\mathcal{Y}\rightarrow\mathcal{Y}$ and $\mathbf{Q}^{\text{ML}}_{k}:\mathcal{Y}\rightarrow\mathcal{V}$ are computed as follows:
$\mathbf{Q}^{\text{ML}}_{k}$ $:=\sum_{\begin{subarray}{c}{i}=1\\ \lambda_{i}\geq 0\end{subarray}}^{N_{{D}}}\widetilde{\mathbf{Q}}^{\text{ML}}_{k}\mathbf{p}_{i}(\mathbf{p}_{i},\,\cdot\,)_{\mathcal{Y}},\;$
$\mathbf{P}^{\text{ML}}_{k}$ $:=\sum_{\begin{subarray}{c}{i}=1\\ \lambda_{i}\geq 0\end{subarray}}^{N_{{D}}}\widetilde{\mathbf{P}}^{\text{ML}}_{k}\mathbf{p}_{i}(\mathbf{p}_{i},\,\cdot\,)_{\mathcal{Y}}=\sum_{\begin{subarray}{c}{i}=1\\ \lambda_{i}\geq 0\end{subarray}}^{N_{{D}}}\lambda_{{i}}\mathbf{p}_{i}(\mathbf{p}_{i},\,\cdot\,)_{\mathcal{Y}}.$
@@
Finally, the regularized covariances are used to define the multi-level Kalman gain $\mathbf{K}_{k}^{\text{ML}}:=\mathbf{Q}^{\text{ML}}_{k}(\mathbf{P}^{\text{ML}}_{k}+\bm{\Sigma})^{-1}$, which is employed to update the states of the three ensembles. Specifically, for all $n\in\{1,...,N_{P}\}$ and $m\in\{1,...,N_{A}\}$, we update the ensemble members as
@@ 4.2 The MF-EnKF analysis step
$\mathbf{P}^{\text{MF}}_{k}$ $=\mathbf{P}^{P,P}_{k}+\dfrac{1}{4}\left(\mathbf{P}^{C,C}_{k}+\mathbf{P}^{A,A}_{k}\right)-\dfrac{1}{2}\left(\mathbf{P}^{P,C}_{k}+\mathbf{P}^{C,P}_{k}\right){\color[rgb]{0,0,0},}$
@@
The multi-fidelity Kalman gain is then defined as $\mathbf{K}_{k}^{\text{MF}}:=\mathbf{Q}^{\text{MF}}_{k}(\mathbf{P}^{\text{MF}}_{k}+\tfrac{1}{2}\bm{\Sigma})^{-1}$, with $\mathbf{P}^{\text{MF}}_{k}$ and $\mathbf{Q}^{\text{MF}}_{k}$ as in (12). The states of the three ensembles are updated, for all $n\in\{1,...,N_{P}\}$ and $m\in\{1,...,N_{A}\}$, as
@@
$\overline{\omega}_{k|k}^{\text{MF}}:=\overline{\omega}_{k|k}^{P}+\dfrac{1}{2}(\overline{\omega}_{k|k}^{A}-\overline{\omega}_{k|k}^{{C}})\;\;\mbox{where}\;\;\overline{\omega}_{k|k}^{\Diamond}=\dfrac{1}{N_{\Diamond}}\textstyle\sum_{n=1}^{N_{\Diamond}}\tilde{\omega}_{k|k}^{{\Diamond,n}},\;\;\Diamond\in\{P,C,A\}.$
@@ 4.3 The prediction step: an inflation-deflation adaptive reduced basis approach
At each iteration of the EnKF algorithm, we construct two subspaces of $\mathcal{V}$: the first one $\mathcal{V}_{\varepsilon}^{k}$, called the inflated space, is the reduced basis space; the second one $\mathcal{W}_{\varepsilon}^{k}$, called the deflated space, is the approximation space that transfers information from one assimilation window to the subsequent one. Before starting the data assimilation process, along with the initial ensembles, it is necessary to specify the starting deflated space, $\mathcal{W}_{\varepsilon}^{0}$, to incorporate any potential prior knowledge regarding the model’s reducibility. In the absence of more information, one can assume $\mathcal{W}_{\varepsilon}^{0}\equiv\mathcal{V}$. Next, at the $k$th iteration of the data assimilation algorithm, the prediction step starts with the forecast of the post-analysis principal ensemble, $E_{k|k}^{P}$, over the assimilation window $\mathcal{I}_{k}$. Specifically, the high-fidelity model ${\color[rgb]{0,0,0}\Psi(\,\cdot\,,t_{k},t_{k+1})}$ is employed to update the states in the principal ensemble
$\omega_{k+1|k}^{P,n}={\color[rgb]{0,0,0}\Psi(\omega_{k|k}^{P,n},t_{k},t_{k+1})},\quad\forall\,n\in\{1,...,N_{P}\}.$ (15)
@@
$\mathcal{V}_{\varepsilon}^{k+1}:=\text{span}\{\text{POD}(\Xi_{\Delta,k}^{\text{HF}},\varepsilon^{\Diamond}_{k}/2),\mathcal{W}_{\varepsilon}^{k}\},$ (16)
@@
$\varepsilon^{\text{MF}}_{k}=\,\dfrac{2\varepsilon_{r}}{N_{P}-1}\sum_{n=1}^{N_{P}}\Big{(}\|\omega_{k|k}^{P,n}-\overline{\omega}_{k|k}^{P}\|_{\mathcal{V}}^{2}-(\omega_{k|k}^{P,n}-\overline{\omega}_{k|k}^{P},\omega_{k|k}^{C,n}-\overline{\omega}_{k|k}^{C})_{\mathcal{V}}$ (18)
$+\dfrac{1}{4}\|\omega_{k|k}^{C,n}-\overline{\omega}_{k|k}^{C}\|_{\mathcal{V}}^{2}\Big{)}+\dfrac{1}{2}\dfrac{\varepsilon_{r}}{N_{A}-1}\sum_{n=1}^{N_{A}}\|\omega_{k|k}^{A,m}-\overline{\omega}_{k|k}^{A}\|_{\mathcal{V}}^{2}.$
@@
The resulting approximation space $\mathcal{V}_{\varepsilon}^{k+1}$ of size $N_{\varepsilon}^{k+1}$ is used to construct the surrogate model ${\color[rgb]{0,0,0}\Psi^{k+1}_{\varepsilon}(\,\cdot\,,t_{k},t_{k+1})}$, following the procedure of Section 3. The updated model is then employed to advance the states in the control and ancillary ensembles, namely
$\omega_{k+1|k}^{C,n}$ $={\color[rgb]{0,0,0}\Psi^{k+1}_{\varepsilon}(\Pi_{\mathcal{V}_{\varepsilon}^{k+1}}\omega_{k|k}^{P,n},t_{k},t_{k+1})},$ $\quad\forall\,n\in\{1,...,N_{P}\},$ (19)
$\omega_{k+1|k}^{A,m}$ $={\color[rgb]{0,0,0}\Psi^{k+1}_{\varepsilon}(\Pi_{\mathcal{V}_{\varepsilon}^{k+1}}\omega_{k|k}^{A,m},t_{k},t_{k+1})},$ $\quad\forall\,m\in\{1,...,N_{A}\}.$
@@
respectively. These are used to deflate $\mathcal{V}_{\varepsilon}^{k+1}$, retaining only the modes actively employed by the surrogate model: if $\Xi_{\Delta}^{\text{LF}}:=\mathcal{E}_{k}^{C}\cup\mathcal{E}_{k}^{A}$, the updated deflated space is
$\mathcal{W}_{\varepsilon}^{k+1}:=\text{POD}(\Xi_{\Delta}^{\text{LF}},\varepsilon^{\Diamond}_{k}/2),$ (20)
@@
The rationale behind the choice of the POD tolerance in Equation 16 and Equation 20 is to build an approximation space with a prescribed tolerance $\varepsilon^{\Diamond}_{k}$ by means of two intermediate approximation spaces whose error contributions could be orthogonal. Intuitively, by ensuring that each of the two spaces introduces an average error equal to half the tolerance $\varepsilon^{\Diamond}_{k}$, we aim at obtaining a final model error that is lower than or equal to $\varepsilon^{\Diamond}_{k}$. This is, however, only an empirical choice. In fact, the approximation error associated with the space constructed in the deflation phase refers to a different time window from the one in which the inflated model is designed to operate. This involves an additional error contribution that is difficult to estimate.